# Affine Chabauty I

### v1

@@ 3.4. Reduction types for more general models
Let $S$ be a finite set of primes of $\mathcal{O}_{K}$. Let $Y=X\smallsetminus D$ as before and let $\mathcal{X}$ be a model of $X$ over $\mathcal{O}_{K}$ [Liu06, Definition 10.1.1] which is not necessarily regular and not necessarily $D$-transversal over primes in $S$. Let $\operatorname{\mathcal{D}}$ be the closure of $D$ in $\mathcal{X}$ and set $\mathcal{Y}=\mathcal{X}\smallsetminus\operatorname{\mathcal{D}}$. We can still partition the set of $S$-integral points $\mathcal{Y}(\mathcal{O}_{K,S})$ into subsets on which the images under an Abel–Jacobi map $Y(K)\to J_{Y}(K)$ are strongly constrained, but we have to use reduction types on a suitably desingularised model. Let $\pi\colon\mathcal{X}^{\prime}\to\mathcal{X}$ be a morphism of models over $\mathcal{O}_{K}$ where $\mathcal{X}^{\prime}$ is regular and $D$-transversal over primes in $S$. Such a desingularisation can be obtained as follows: by repeatedly blowing up in the singular locus and normalising, one obtains a regular model [Liu06, Cor. 8.3.51], and then one can use ˜2.19. Let $\operatorname{\mathcal{D}}^{\prime}$ be the closure of $D$ in $\mathcal{X}^{\prime}$ and set $\mathcal{Y}^{\prime}=\mathcal{X}^{\prime}\smallsetminus\operatorname{\mathcal{D}}^{\prime}$. We have the inclusion $\mathcal{Y}(\mathcal{O}_{K,S})\subseteq\mathcal{Y}^{\prime}(\mathcal{O}_{K,S})$, which can be strict, for example if $\pi$ contracts a vertical component onto a point lying on $\operatorname{\mathcal{D}}$. But by allowing only certain reduction types, we can find $\mathcal{Y}(\mathcal{O}_{K,S})$ inside $\mathcal{Y}^{\prime}(\mathcal{O}_{K,S})$ and thereby reduce the computation of $S$-integral points on $\mathcal{Y}$ to the case of a model which is regular and $D$-transversal over primes in $S$.
@@
  • for primes $\mathfrak{q}\not\in S$: an irreducible (equivalently: connected) component of ${\mathcal{X}^{\prime}_{\mathfrak{q}}}^{\mathrm{sm}}\smallsetminus\pi^{-1}(\operatorname{\mathcal{D}})$ that contains a smooth $\operatorname{\mathbb{F}}_{\mathfrak{q}}$-point;
  • for primes $\mathfrak{q}\in S$: a component of ${\mathcal{X}^{\prime}_{\mathfrak{q}}}^{\mathrm{sm}}\smallsetminus\operatorname{\mathcal{D}}^{\prime}$ that contains a smooth $\operatorname{\mathbb{F}}_{\mathfrak{q}}$-point, or a point in $({\mathcal{X}^{\prime}_{\mathfrak{q}}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}}^{\prime})(\operatorname{\mathbb{F}}_{\mathfrak{q}})$.
@@ 4. $p$-adic integrals of logarithmic differentials
Let $K$ be a complete subfield of $\mathbb{C}_{p}$ (such as $\mathbb{Q}_{p}$ or a finite extension thereof), let $X$ be a smooth proper curve of genus $g$ over $K$, let $D\neq\emptyset$ be a finite set of closed points of $X$ and set $Y=X\smallsetminus D$. Denote by $n\coloneqq D(\overline{K})$ the number of geometric cusps. A logarithmic differential on $(X,D)$ is a meromorphic differential on $X$ which is everywhere regular except for possibly simple poles at the points of $D$. The $K$-vector space of logarithmic differentials on $(X,D)$ is $\mathrm{H}^{0}(X,\Omega^{1}(D))$. By the Riemann–Roch theorem, its dimension is given by
@@ 5. Construction of Chabauty functions
We now prove ˜A over number fields as well as several variants. Let $K$ be a number field. Let $X$ be a smooth proper curve over $K$, let $D\neq\emptyset$ be a finite set of closed points of $X$ and set $Y=X\smallsetminus D$. Let $n=D(\overline{K})$. Let $S$ be a finite set of primes of $\mathcal{O}_{K}$ and let $\mathcal{X}$ be a regular model of $X$ over $\mathcal{O}_{K}$ that is $D$-transversal over primes in $S$. Fix a base point $P_{0}\in Y(K)$ and an $S$-integral reduction type $\Sigma=(\Sigma_{\mathfrak{q}})_{\mathfrak{q}}$ for $(\mathcal{X},\operatorname{\mathcal{D}})$ (see ˜3.8). We are interested in the set $\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$ consisting of all $S$-integral points of reduction type $\Sigma$.
@@ 5.1. Existence of Chabauty functions
Then there exist a non-zero log differential $0\neq\omega\in\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$ and a constant $c\in K_{\mathfrak{p}}$ such that the function $\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})\to K_{\mathfrak{p}}$ given by
@@
vanishes on $\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$.
Hypothesis (5.1) specialises to (1.2) for $K=\mathbb{Q}$, so ˜A is a special case of ˜5.1. Note that the function $\rho\colon\mathcal{Y}(\mathbb{Z}_{p})\to\mathbb{Q}_{p}$ in (5.2) has only finitely many zeros by §4.2.
@@
If $D$ is just a single $K$-rational point, Condition (5.1) becomes $r+\#C(\Sigma)<g$, and already the weaker condition $r<g$ is sufficient for the classical Chabauty method to apply, yielding a non-zero holomorphic differential $0\neq\omega\in\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1})$ such that $\int_{P_{0}}^{P}\omega=0$ for all rational points $P\in X(K)$. Hence, from now on assume $n>1$. In Section˜3.3 we constructed the Selmer set $\operatorname{Sel}(P_{0},\Sigma)\subseteq J_{Y}(K)$ with the property that the Abel–Jacobi map $\operatorname{AJ}_{P_{0}}\colon\mathcal{Y}(\mathcal{O}_{K,S})\to J_{Y}(K)$ maps all points of reduction type $\Sigma$ into $\operatorname{Sel}(P_{0},\Sigma)$ (see ˜3.13). Consider the following diagram:
$\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$$\mathcal{Y}(\mathcal{O}_{K,S})$$\mathcal{Y}(\mathcal{O}_{K,\mathfrak{p}})$$\operatorname{Sel}(P_{0},\Sigma)$$J_{Y}(K)$$J_{Y}(K_{\mathfrak{p}})$$\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))^{\vee}$$\scriptstyle{\operatorname{AJ}_{P_{0}}}$$\scriptstyle{\operatorname{AJ}_{P_{0}}}$$\scriptstyle{\operatorname{AJ}_{P_{0}}}$$\scriptstyle{\int_{P_{0}}}$$\scriptstyle{\log_{J_{Y}}}$ (5.3)
The set $\operatorname{Sel}(P_{0},\Sigma)$ is the preimage under $\sigma\colon J_{Y}(K)\to V_{D}$ of a translate of a certain subgroup $U\subseteq V_{D}$. Let $W\coloneqq\sigma^{-1}(U)\subseteq J_{Y}(K)$, so that $\operatorname{Sel}(P_{0},\Sigma)$ is either empty or a translate of $W$. When the rank of $W$ is strictly less than $\dim_{K_{\mathfrak{p}}}\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$, then the image of $W$ in $\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))^{\vee}$ under the logarithm map is contained in a proper $K_{\mathfrak{p}}$-linear subspace, so there exists a non-zero $0\neq\omega\in\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$ with $\log_{J_{Y}}(w)(\omega)=0$ for all $w\in W$. Then, since $\operatorname{Sel}(P_{0},\Sigma)$ is a translate of $W$ (or empty), the map $s\mapsto\log_{J_{Y}}(s)(\omega)$ is constant on $\operatorname{Sel}(P_{0},\Sigma)$. Pulling this back under the diagonal map in the diagram (5.3) shows that the map $P\mapsto\int_{P_{0}}^{P}\omega$ is constant on $\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$. By ˜3.14, the condition that the rank of $W$ be strictly less than the dimension of $\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$ reads
@@
The proof above shows that the differential $\omega$, the constant $c$ and hence the function $\rho$ in (5.2) only depend on $\operatorname{Sel}(P_{0},\Sigma)$. That is, if $\operatorname{Sel}(P_{0},\Sigma)=\operatorname{Sel}(P_{0},\Sigma^{\prime})$ for two reduction types $\Sigma$ and $\Sigma^{\prime}$, then $\rho$ in fact vanishes on $\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}\cup\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma^{\prime}}$. The equality $\operatorname{Sel}(P_{0},\Sigma)=\operatorname{Sel}(P_{0},\Sigma^{\prime})$ is of course guaranteed if $\mathfrak{S}_{\mathfrak{q}}(P_{0},\Sigma)=\mathfrak{S}_{\mathfrak{q}}(P_{0},\Sigma^{\prime})$ for all $\mathfrak{q}$.
5.2. Reduction types with fixed cuspidal part
We can say something more precise about how the Chabauty function obtained in ˜5.1 depends on the reduction type. Namely, the differential $\omega$ can be chosen uniformly for all reduction types having the same “cuspidal part”. For a subset $S_{0}\subseteq S$ and a tuple $\Sigma^{c}\in\prod_{\ell\in S_{0}}\mathcal{C}_{\ell}^{1}$, we say that an $S$-integral reduction type $\Sigma=(\Sigma_{\ell})_{\ell}$ has cuspidal part $\Sigma^{c}$ if the restriction of $\Sigma$ to $S_{0}$ is given by $\Sigma^{c}$ and $\Sigma_{\ell}\in\mathcal{C}_{\ell}^{0}$ for $\ell\not\in S_{0}$. In other words, if $\Sigma_{\ell}$ is cuspidal, then $\Sigma_{\ell}=\Sigma^{c}_{\ell}$. Note that in the case $S=\emptyset$, all ($S$-)integral reduction types automatically have the same cuspidal part.
Theorem 5.3.
Let $S_{0}$ be a set of primes in $S$ and fix $\Sigma^{c}\in\prod_{\ell\in S_{0}}\mathcal{C}^{1}_{\ell}$. Assume that the Affine Chabauty Condition
$r+\#S_{0}+([K:\mathbb{Q}]-1)n<g+\operatorname{\mathrm{rk}}\mathcal{O}_{K}^{\times}+\#|D|+n_{2}(D)-1$ (5.4)
is satisfied. Then there exists a non-zero log differential $0\neq\omega\in\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$ such that for all $S$-integral reduction types $\Sigma$ with cuspidal part $\Sigma^{c}$, the function $\rho\colon\mathcal{Y}(\mathcal{O}_{K,\mathfrak{p}})\to K_{\mathfrak{p}}$, $\rho(P)=\int_{P_{0}}^{P}\omega$ is constant on $\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$.
Proof.
Let $U\subseteq V_{D}$ be the subgroup generated by the elements $\Sigma_{\ell}]$ for $\ell\in S_{0}$ and let $W\coloneqq\sigma^{-1}(U)$ be its preimage under the $D$-intersection map $\sigma\colon J_{Y}(K)\to V_{D}$. Then for any reduction type $\Sigma$ with cuspidal part $\Sigma^{c}$, the set $\mathfrak{S}(P_{0},\Sigma)$ defined in ˜3.9 is a translate of $U$, and consequently the Selmer set $\operatorname{Sel}(P_{0},\Sigma)=\sigma^{-1}(\mathfrak{S}(P_{0},\Sigma))$ is empty or a translate of $W$. Hence the proof of ˜5.1 shows the result. ∎
@@ 5.3. Affine Restriction of Scalars Chabauty
Instead of locating $\mathcal{Y}(\mathcal{O}_{K,S})$ inside $\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})$ for some suitable prime $\mathfrak{p}$ of $\mathcal{O}_{K}$ as in Section˜5.1, it is advantageous to consider all primes $\mathfrak{p}$ of $\mathcal{O}_{K}$ lying over a fixed rational prime $p$ at once and try to locate $\mathcal{Y}(\mathcal{O}_{K,S})$ inside $\prod_{\mathfrak{p}|p}\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})$. This method is known as Restriction of Scalars Chabauty. It appeared first in [Sik13] in the context of rational points on projective curves.
@@
If $D$ consists just of a single $K$-rational point, Condition (5.6) is equivalent to $r+\#C(\Sigma)\leq d(g-1)$, and already the weaker condition $r\leq d(g-1)$ is enough to employ the classical Restriction of Scalars Chabauty method for the $K$-rational points on $X$. So we can assume that $n>1$. Consider the diagram (5.5). The Selmer set $\operatorname{Sel}(P_{0},\Sigma)$ constructed in Section˜3.3 is the preimage under $\sigma\colon J_{Y}(K)\to V_{D}$ of a translate of a certain subgroup $U\subseteq V_{D}$. Let $W\coloneqq\sigma^{-1}(U)\subseteq J_{Y}(K)$, so that $\operatorname{Sel}(P_{0},\Sigma)$ is either empty or a translate of $W$. Let $d^{\prime}$ be the codimension of the image of $W\otimes_{\mathbb{Z}}\mathbb{Q}_{p}$ in the $\mathbb{Q}_{p}$-vector space $\prod_{\mathfrak{p}|p}\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))^{\vee}$ along the bottom row of Diagram (5.5). Then there exist $d^{\prime}$ independent $\mathbb{Q}_{p}$-linear maps
@@
When applying ˜5.4 it may be convenient to choose a prime $p$ that is totally split in $K$, so that $K_{\mathfrak{p}}=\mathbb{Q}_{p}$ for all $\mathfrak{p}|p$. Then the integrals occurring in the Chabauty functions (5.7) are all defined over $\mathbb{Q}_{p}$ and the trace map can be omitted. Most implementations of Coleman integration in computer algebra systems are currently restricted to $\mathbb{Q}_{p}$.
@@
In the case that $X:y^{2}=f(x)$ with $f\in\mathcal{O}_{K}[x]$ monic, squarefree and of degree $2g+2$, condition (5.1) specialises to $\operatorname{\mathrm{rk}}J(K)\leq dg-\operatorname{\mathrm{rk}}\mathcal{O}_{K}^{\times}$. In this setting, functions $\prod_{\mathfrak{p}|p}\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})\to\mathbb{Q}_{p}$ vanishing on $\mathcal{Y}(\mathcal{O}_{K})$ are constructed in terms of $p$-adic heights in [GM25, Theorem 5.1].

### v2

@@ 3.4. Reduction types for more general models
Let $S$ be a finite set of primes of $\mathcal{O}_{K}$. Let $Y=X\smallsetminus D$ as before and let $\mathcal{X}$ be a model of $X$ over $\mathcal{O}_{K}$ [Liu06, Definition 10.1.1] which is not necessarily regular and not necessarily $D$-transversal over primes in $S$. Let $\operatorname{\mathcal{D}}$ be the closure of $D$ in $\mathcal{X}$ and set $\mathcal{Y}=\mathcal{X}\smallsetminus\operatorname{\mathcal{D}}$. We can still partition the set of $S$-integral points $\mathcal{Y}(\mathcal{O}_{K,S})$ into subsets whose images under an Abel–Jacobi map $Y(K)\to J_{Y}(K)$ are strongly constrained, but we have to use reduction types on a suitably desingularised model. Let $\pi\colon\mathcal{X}^{\prime}\to\mathcal{X}$ be a morphism of models over $\mathcal{O}_{K}$ where $\mathcal{X}^{\prime}$ is regular and $D$-transversal over primes in $S$. Such a desingularisation can be obtained as follows: by repeatedly blowing up in the singular locus and normalising, one obtains a regular model [Liu06, Cor. 8.3.51], and then one can use ˜2.19. Let $\operatorname{\mathcal{D}}^{\prime}$ be the closure of $D$ in $\mathcal{X}^{\prime}$ and set $\mathcal{Y}^{\prime}=\mathcal{X}^{\prime}\smallsetminus\operatorname{\mathcal{D}}^{\prime}$. We have the inclusion $\mathcal{Y}(\mathcal{O}_{K,S})\subseteq\mathcal{Y}^{\prime}(\mathcal{O}_{K,S})$, which can be strict, for example if $\pi$ contracts a vertical component onto a point lying on $\operatorname{\mathcal{D}}$. But by allowing only certain reduction types, we can find $\mathcal{Y}(\mathcal{O}_{K,S})$ inside $\mathcal{Y}^{\prime}(\mathcal{O}_{K,S})$ and thereby reduce the computation of $S$-integral points on $\mathcal{Y}$ to the case of a model which is regular and $D$-transversal over primes in $S$.
@@
  • for primes $\mathfrak{q}\not\in S$: an irreducible (equivalently: connected) component of ${\mathcal{X}^{\prime}_{\mathfrak{q}}}^{\mathrm{sm}}\smallsetminus\pi^{-1}(\operatorname{\mathcal{D}})$ containing an $\operatorname{\mathbb{F}}_{\mathfrak{q}}$-point;
  • for primes $\mathfrak{q}\in S$: either a component of ${\mathcal{X}^{\prime}_{\mathfrak{q}}}^{\mathrm{sm}}\smallsetminus\operatorname{\mathcal{D}}^{\prime}$ containing an $\operatorname{\mathbb{F}}_{\mathfrak{q}}$-point, or a point in $({\mathcal{X}^{\prime}_{\mathfrak{q}}}^{\mathrm{sm}}\cap\operatorname{\mathcal{D}}^{\prime})(\operatorname{\mathbb{F}}_{\mathfrak{q}})$.
@@ 4. $p$-adic integrals of logarithmic differentials
Let $K$ be a complete subfield of $\mathbb{C}_{p}$ (such as $\mathbb{Q}_{p}$ or a finite extension thereof), let $X$ be a smooth projective curve of genus $g$ over $K$, let $D\neq\emptyset$ be a finite set of closed points of $X$ and set $Y=X\smallsetminus D$. Denote by $n\coloneqq D(\overline{K})$ the number of geometric cusps. A logarithmic differential on $(X,D)$ is a meromorphic differential on $X$ which is everywhere regular except for possibly simple poles at the points of $D$. The $K$-vector space of logarithmic differentials on $(X,D)$ is $\mathrm{H}^{0}(X,\Omega^{1}(D))$. By the Riemann–Roch theorem, its dimension is given by
@@ 5. Construction of Chabauty functions
We now prove ˜A over number fields and describe the affine variant of Restriction of Scalars Chabauty. Let $K$ be a number field. Let $X$ be a smooth projective curve over $K$, let $D\neq\emptyset$ be a finite set of closed points of $X$ and set $Y=X\smallsetminus D$. Let $n=\#D(\overline{K})$. Let $S$ be a finite set of primes of $\mathcal{O}_{K}$ and let $\mathcal{X}$ be a regular model of $X$ over $\mathcal{O}_{K}$ that is $D$-transversal over primes in $S$. Fix a base point $P_{0}\in Y(K)$ and an $S$-integral reduction type $\Sigma=(\Sigma_{\mathfrak{q}})_{\mathfrak{q}}$ for $(\mathcal{X},\operatorname{\mathcal{D}})$ (see ˜3.8). We are interested in the set $\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$ consisting of all $S$-integral points of reduction type $\Sigma$. Recall that $C(\Sigma)$ denotes the set of primes $\mathfrak{q}\in S$ for which $\Sigma_{\mathfrak{q}}$ is cuspidal, and $\Sigma^{{\mathrm{csp}}}\coloneqq(\Sigma_{\mathfrak{q}})_{\mathfrak{q}\in C(\Sigma)}$ is called the cuspidal part of $\Sigma$ (˜3.10).
@@ 5.1. Existence of Chabauty functions
Then there exist a non-zero log differential $0\neq\omega\in\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$ and a constant $c\in K_{\mathfrak{p}}$, both depending on $\Sigma$, such that the function $\rho\colon\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})\to K_{\mathfrak{p}}$ given by
@@
vanishes on $\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$. The differential $\omega$ can be chosen uniformly for all reduction types having the same cuspidal part.
Hypothesis (5.1) specialises to (1.2) for $K=\mathbb{Q}$, so ˜A is a special case of ˜5.1. Note that the function $\rho\colon\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})\to K_{\mathfrak{p}}$ in (5.2) has only finitely many zeros by §4.2.
@@
If $D$ is just a single $K$-rational point, Condition (5.1) becomes $r+\#C(\Sigma)<g$, and already the weaker condition $r<g$ is sufficient for the classical Chabauty method to apply, yielding a non-zero holomorphic differential $0\neq\omega\in\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1})$ such that $\int_{P_{0}}^{P}\omega=0$ for all rational points $P\in X(K)$. Hence, from now on assume $n>1$. In Section˜3.3 we constructed the Selmer set $\operatorname{Sel}(P_{0},\Sigma)\subseteq J_{Y}(K)$ with the property that the Abel–Jacobi map $\operatorname{AJ}_{P_{0}}\colon\mathcal{Y}(\mathcal{O}_{K,S})\to J_{Y}(K)$ maps all points of reduction type $\Sigma$ into $\operatorname{Sel}(P_{0},\Sigma)$ (see ˜3.13). The set $\operatorname{Sel}(P_{0},\Sigma)$ is the preimage under $\sigma\colon J_{Y}(K)\to V_{D}$ of a translate of a certain subgroup $U=U(\Sigma^{{\mathrm{csp}}})\subseteq V_{D}$. Let $W\coloneqq\sigma^{-1}(U)\subseteq J_{Y}(K)$, so that $\operatorname{Sel}(P_{0},\Sigma)$ is either empty or a translate $s_{0}+W$ of $W$. Consider the following diagram:
$\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$$\mathcal{Y}(\mathcal{O}_{K,S})$$\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})$$s_{0}+W$$\operatorname{Sel}(P_{0},\Sigma)$$J_{Y}(K)$$J_{Y}(K_{\mathfrak{p}})$$\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))^{\vee}$$\scriptstyle{\operatorname{AJ}_{P_{0}}}$$\scriptstyle{\operatorname{AJ}_{P_{0}}}$$\scriptstyle{\operatorname{AJ}_{P_{0}}}$$\scriptstyle{\int_{P_{0}}}$$\scriptstyle{\log_{J_{Y}}}$ (5.3)
When $\operatorname{\mathrm{rk}}_{\mathbb{Z}}W$ is strictly less than $\dim_{K_{\mathfrak{p}}}\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$, then the image of $W$ in $\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))^{\vee}$ under the logarithm map is contained in a proper $K_{\mathfrak{p}}$-linear subspace, so there exists a non-zero $0\neq\omega\in\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$ with $\langle\log_{J_{Y}}(w),\omega\rangle=0$ for all $w\in W$. Then, since $\operatorname{Sel}(P_{0},\Sigma)$ is a translate of $W$ (or empty), the map $s\mapsto\langle\log_{J_{Y}}(s),\omega\rangle$ is constant on $\operatorname{Sel}(P_{0},\Sigma)$. Pulling this back under the diagonal map in the diagram (5.3) shows that the map $P\mapsto\int_{P_{0}}^{P}\omega$ is constant on $\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}$. By ˜3.14, the condition that the rank of $W$ be strictly less than the dimension of $\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))$ reads
@@
The proof above shows that the differential $\omega$, the constant $c$ and hence the function $\rho$ in (5.2) depend on $\Sigma$ only through $\operatorname{Sel}(P_{0},\Sigma)$. That is, if $\operatorname{Sel}(P_{0},\Sigma)=\operatorname{Sel}(P_{0},\Sigma^{\prime})$ for two reduction types $\Sigma$ and $\Sigma^{\prime}$, then we get a single Chabauty function $\rho$ that vanishes on $\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma}\cup\mathcal{Y}(\mathcal{O}_{K,S})_{\Sigma^{\prime}}$. The equality $\operatorname{Sel}(P_{0},\Sigma)=\operatorname{Sel}(P_{0},\Sigma^{\prime})$ is guaranteed if $\mathfrak{S}_{\mathfrak{q}}(P_{0},\Sigma)=\mathfrak{S}_{\mathfrak{q}}(P_{0},\Sigma^{\prime})$ for all $\mathfrak{q}$.
@@ 5.2. Affine Restriction of Scalars Chabauty
Instead of locating $\mathcal{Y}(\mathcal{O}_{K,S})$ inside $\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})$ for some suitable prime $\mathfrak{p}$ of $\mathcal{O}_{K}$ as in Section˜5.1, it can be advantageous to consider all primes $\mathfrak{p}$ of $\mathcal{O}_{K}$ lying over a fixed rational prime $p$ at once and try to locate $\mathcal{Y}(\mathcal{O}_{K,S})$ inside $\prod_{\mathfrak{p}|p}\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})$. This method is known as Restriction of Scalars Chabauty. It appeared first in [Sik13] in the context of rational points on projective curves.
@@
If $D$ consists just of a single $K$-rational point, Condition (5.5) is equivalent to $r+\#C(\Sigma)\leq d(g-1)$, and already the weaker condition $r\leq d(g-1)$ is enough to employ the classical Restriction of Scalars Chabauty method for the $K$-rational points on $X$. So we can assume that $n>1$. Consider the diagram (5.4). The Selmer set $\operatorname{Sel}(P_{0},\Sigma)$ constructed in Section˜3.3 is the preimage under $\sigma\colon J_{Y}(K)\to V_{D}$ of a translate of a certain subgroup $U=U(\Sigma^{{\mathrm{csp}}})\subseteq V_{D}$. Let $W\coloneqq\sigma^{-1}(U)\subseteq J_{Y}(K)$, so that $\operatorname{Sel}(P_{0},\Sigma)$ is either empty or a translate of $W$. Let $d^{\prime}$ be the codimension of the image of $W\otimes_{\mathbb{Z}}\mathbb{Q}_{p}$ in the $\mathbb{Q}_{p}$-vector space $\prod_{\mathfrak{p}|p}\mathrm{H}^{0}(X_{K_{\mathfrak{p}}},\Omega^{1}(D))^{\vee}$ along the bottom row of Diagram (5.4). Then there exist $d^{\prime}$ independent $\mathbb{Q}_{p}$-linear maps
@@
When applying ˜5.3 it may be convenient to choose a prime $p$ that is totally split in $K$, so that $K_{\mathfrak{p}}=\mathbb{Q}_{p}$ for all $\mathfrak{p}|p$. Then the integrals occurring in the Chabauty functions (5.6) are all defined over $\mathbb{Q}_{p}$ and the trace map can be omitted. Most implementations of Coleman integration in computer algebra systems are currently restricted to $\mathbb{Q}_{p}$.
@@
In the case that $\mathcal{Y}\colon y^{2}=f(x)$ with $f\in\mathcal{O}_{K}[x]$ monic, squarefree and of degree $2g+2$, condition (5.5) with $S=\emptyset$ specialises to $\operatorname{\mathrm{rk}}J(K)\leq dg-\operatorname{\mathrm{rk}}\mathcal{O}_{K}^{\times}$. In this setting, functions $\prod_{\mathfrak{p}|p}\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})\to\mathbb{Q}_{p}$ vanishing on $\mathcal{Y}(\mathcal{O}_{K})$ are constructed in terms of $p$-adic heights in [GM25, Theorem 5.1].